# Dissecting CLIP: Decomposition with a Schur Complement-based Approach

### v2

@@ 5 Numerical Results
Figure 3: Evaluated $\mathrm{SCE_{I}}$ scores with fixed text prompt distribution about doctors in various settings. Interpolated surface plot in the bottom right visualises the change in diversity (z-axis) when doctors have a varying diversity of age and gender features. (x and y axis)
@@
We evaluated our proposed Schur Complement-based approach for CLIP decomposition across various scenarios using both real and synthetic datasets. Our experimental results are reported for the standard Cosine Similarity and Gaussian kernel functions. For the Gaussian kernel function, we use random Fourier features [26] with a random Fourier dimension $r=2000$ to embed the kernel using a finite feature dimension. In the diversity evaluation, we report our defined SCE scores as well as the Vendi score [6], RKE[18], Recall[4] and Coverage[5] for unconditional diversity assessment without taking the text data into account. Note that Recall and Coverage may only be performed in the presence of a valid representative reference dataset. The clustering experiments were performed using Kernel PCA with the Gaussian kernel, highlighting the top clusters that align with the top eigenvector directions in the data.
Measuring Text-to-Image Model Diversity. To highlight the strengths of SCE-based diversity evaluation, we conducted numerical experiments comparing unconditional and conditional diversity assessments. Figure 3 presents an experiment with the FLUX.1-dev [27] model, where individuals are generated based on an ambiguous prompt. We fixed the text prompt distribution and conditioned the model to generate doctors of varying gender and age. The results show that when the model is restricted to producing only "senior male doctors," the $\mathrm{SCE_{I}}$ score is the lowest. Conversely, when the model generates a broader range of age groups and genders, the diversity score increases. We also provide an interpolated surface plot illustrating the intermediate mixing results. A diversity level of 0 indicates a strong bias toward a single characteristic (e.g., all images depict males), whereas a score of 1 signifies an equal representation of all characteristics (e.g., an equal number of male and female images).
Figure 4: Evaluated $\mathrm{SCE_{I}}$ scores with fixed text prompt distribution. $\mathrm{SCE_{I}}$ changes to reflect the diversity contributed by the varying level of information in text prompts.
Figure 4 illustrates the inverse scenario: the image distribution is fixed while the accompanying text prompts vary in their specificity. We generated four sets of "junior male doctor" images with different seeds, while prompts were assigned four levels of detail—fully specified (age and gender), partially specified (age only or gender only), or unspecified (no details). We measured $\mathrm{SCE_{I}}$, Vendi, RKE, Coverage and Recall scores across four scenarios. Reference set for R/C is "doctor" class in IdenProf dataset [28]. Vendi, RKE, Recall and Coverage scores remain relatively unchanged, whereas $\mathrm{SCE_{I}}$ distinguishes between these cases. When the description is incomplete, $\mathrm{SCE_{I}}$ is lowest due to the strong presence of "junior male doctors" across all images, indicating bias. However, when prompts explicitly state that all images should depict "junior male doctors," diversity increases and is evaluated based on auxiliary features such as background details rather than age or gender. When partial information is given, $\mathrm{SCE_{I}}$ falls between the fully specified and unspecified cases. These results highlight the need for differentiation between model-driven and prompt-driven diversity.
To further quantify our findings, in Figure 5 we conducted an experiment on SDXL generated cat images of different breeds. We generated 1,000 images per class and compared $\mathrm{SCE_{I}}$ with other diversity metrics. When prompts did not specify cat breeds, the diversity captured by $\mathrm{SCE_{I}}$ aligns with unconditional Vendi score. Conversely, when breed information was included in the prompts, diversity mainly stemmed from the text prompts rather than the image generator, leading to a non-increasing $\mathrm{SCE_{I}}$ score. In both scenarios, Vendi, Coverage, and Recall scores remained unchanged. The reference set for Recall and Coverage consisted of ImageNet [29] ’cat’ samples. Additional experiments on other cases are discussed in the Appendix.
Figure 5: Evaluated $\mathrm{SCE_{I}}$, Vendi, Recall and Coverage scores with Gaussian Kernel on different cat breeds dataset.
Figure 6: t-SNE plot of animals with objects dataset.
The presented experimental results highlight the importance of disentangling the sources of diversity in text-to-image generation—specifically, the diversity induced by the text prompts versus that originating from the generative model itself. Our findings reveal that when the distribution of prompts or the resulting images is varied, a reliable evaluation metric must be sensitive to both aspects, adapting accordingly to reflect shifts in either distribution. In contrast, existing metrics primarily assess diversity based solely on image features, overlooking the influence of prompt variability. This limitation reduces their effectiveness in scenarios where prompt semantics play a central role, such as in text-to-image generation. These observations motivate the development of a new metric that jointly captures the interplay between prompt diversity and image diversity, enabling more faithful and nuanced evaluation of generative models in this setting.
(a) Classification distribution before and after CLIP correction on SDXL generated images of animals with objects in the background. The correction is performed via text descriptions of concepts to be removed.
(b) Classification distribution before and after CLIP correction on gpt4o-mini captioned text descriptions about animals with objects in the background. The correction is performed using provided images that encapsulate the concept to be removed.
Figure 7: Classification distribution before and after correction in text-to-image generation and image captioning (image-to-text) tasks
@@
Figure 8: Clusters of DALL-E 3 [7] generated images of animals with fruits. The yellow column shows Kernel-PCA (KPCA) clusters using CLIP; the green column shows KPCA clusters with the application the proposed Schur-Complement-based method to remove "fruit" direction from CLIP embedding; and the blue column shows clusters with the "animal" direction removed from CLIP embedding.
Figure 10: Effect of removing encoded "cassette player" text on top of ImageNet samples.
@@
CLIP Decomposition. To illustrate the effect of Schur-complement-based decomposition on the embedding space, we first generate an "animals with objects" dataset with SDXL, where each sample contains one object from ‘boat’, ‘skyscraper’, ‘basketball’, ‘couch’, ‘umbrella’, ‘sunflowers’ paired with one animal from ‘cat’, ‘dog’, ‘guinea pig’, ‘snake’. In this dataset, diversity increases primarily due to the higher number of subjects rather than feature variations. Using this dataset, we apply t-SNE to visualize the dimensionality-reduced embeddings of CLIP and its corrected variant in Figure 6. The results show that distinct clusters of objects and animals merge when we remove either the "animal" or "object" direction from the embedding, demonstrating the impact of our decomposition. Performing $\mathrm{SCE}$ decomposition causes a change in the embedding space and correlation of features along removed direction greatly decreases.
[…]
Figure 7b further extends our analysis by applying the $\mathrm{SCE}$-based decomposition not only to the text-to-image setting but also to its inverse: the image-to-text, or image captioning, task. The experimental setup mirrors that of Figure 7a, with the key difference being that, instead of generating images for text prompts, we generated captions for the corresponding images. For this task, we used gpt4o-mini as the captioning model. We selected images closely aligned with the concept we aimed to remove. For instance, to eliminate the "cat" direction in text, we used an image of a cat against a white background to better isolate the concept. After applying corrections for "animals" or "objects" in the text prompt, we observed successful decomposition, as reflected in the second column: the corrected CLIP embedding is no longer sensitive to the removed concept. The results closely mirror those observed in the text-to-image case, reinforcing that our decomposition method consistently removes the influence of the specified concept across both textual and visual modalities. This symmetry highlights the robustness of our approach in isolating concept-driven variations, regardless of the generation direction. Our method can be readily extended to other paired tasks, provided that an appropriate embedding model exists to encode the paired modalities.
@@
Additionally, we use the eigenspace of the Schur complement component to interpret the diversity in the generated images that is independent of the text prompt, revealing the unique elements introduced by the generation model in the images. The eigenspace-based interpretation follows the application of the Kernel PCA (KPCA) clustering method and leads to a visualization of the sample clusters due to the input text clusters and the clusters with a feature added by the model to the generated image. Figure 8 shows the resulting KPCA-detected clusters for DALL-E 3 [7] generated images in response to prompts of type "[animal] eating a [fruit]" with 5 different animal and fruit names. The detected clusters of CLIP image embedding (leftmost case) shows clusters with mixed animals and fruits. On the other hand, by considering the prompt of only specifying the fruit, the Schur complement component’s KPCA clusters highlight the animals in each cluster (middle case), and similarly by considering the prompt of only specifying the fruit, the animals are captured by each KPCA cluster of the Schur complement matrix. Further decomposition results and analysis are provided in the Appendix.
Moreover, to test how CLIP correction affects the underlying directions of concepts, we applied CLIPDiffusion [30] framework to edit the image according to different CLIP embeddings. Figure 9 illustrates the setup of the problem, where we edit ’initialization image’ that consists of two subjects: a cat and a basketball. Next we denoise and guide the generation according to three different embeddings: Unchanged CLIP embedding of ’initialization image’, corrected CLIP by a ’cat’ direction or corrected CLIP by a ’basketball’ direction. We show that after removing a concept direction, denoiser is no longer rewarded for generating corresponding concept, which is reflected in denoised images. A basketball after correction resembles a bowl with plants and cat loses all of its features. We also note that in both cases, the other object remains intact. To further showcase these results, we performed the same diffusion on the animals with traffic signs dataset shown on the side of Figure 9. This experiment offers a visual interpretation of how the decomposition process removes specific directional components and what the resulting corrected embedding represents. While the $\mathrm{SCE}$-based decomposition is not intended for direct application in image editing tasks, these results are illustrative of the changes that occur in the embedding space when a particular concept is removed from an image. To further validate our approach, we include a baseline where the correction matrix is set to $\Gamma^{*}=\mathrm{I}$, effectively bypassing the optimization step. In this setting, the resulting embeddings fail to preserve the original semantic directions and deviate significantly from the expected outcomes, underscoring the importance of computing the optimal correction matrix $\Gamma^{*}$ for meaningful concept removal. The additional results are provided in the Appendix.
Typographic Attacks. CLIP’s sensitivity to text within images, as showed in [13], makes it vulnerable to typographic attacks, where overlaying misleading text on an image influences its classification [31]. To investigate this, we constructed a dataset of 10 ImageNet classes, each image engraved with a random class label. We conducted an experiment to evaluate the classification performance of CLIP before and after applying our correction method. Figure 10 presents the distribution of predicted classes for images overlaid with the text "cassette player." Prior to correction, CLIP misclassified nearly 79% of these images as "cassette player," indicating a strong susceptibility to typographic bias. However, after applying the $\mathrm{SCE}$-based correction, the misclassification rate drops dramatically to 12%, demonstrating the effectiveness of our approach in mitigating the effect of a typographic attack in the embedding space.
@@
Furthermore, Figure 11 demonstrates CLIP’s susceptibility, showing that top eigenvector directions capture the engraved text rather than the actual image content. By conditioning on prompts like "text reading ’cassette player’ on top of image" and removing this direction, we re-clustered images based on the corrected CLIP embedding. This adjustment allowed us to identify clusters based on the actual image content rather than the engraved text. We visualize the top modes associated with the leading eigenvectors, revealing that the original CLIP embeddings tend to cluster images based on the overlaid text (e.g., text in the center of the image), rather than the actual image content. In contrast, the $\mathrm{SCE}$-corrected CLIP embeddings cluster images according to their visual content, effectively ignoring the spurious text signal.
@@ 6 Conclusion
In this work, we proposed a Schur Complement-based approach to decompose the kernel covariance matrix of CLIP image embeddings into the sum of text-induced and model-induced kernel covariance components. We demonstrated the application of this Schur Complement-based approach for evaluating the diversity of text-to-image models’ generated data based on the model’s input text. This method extends the application of CLIP embeddings from relevance evaluation to assessing diversity in text-to-image generation models. Additionally, we applied our approach to modify CLIP image embeddings by modifying or canceling the influence of an input text. Our numerical results showcase the potential for canceling text effects and its applications to computer vision tasks. Future research directions include extending the Schur Complement-based approach to other generative AI models, such as text-to-video and language models. Additionally, leveraging this decomposition to fine-tune the CLIP model for enhanced understanding of diverse concepts represents another future direction.
@@ A.1 Proof of Proposition 1
A.2 Conditional Entropy Interpretation of SCE Score
As discussed in the main text, in Equation equation 6, both image component $\Lambda_{I}$ and text component $\Lambda_{T}$ are PSD matrices with unit trace. Furthermore, we have
@@ A.2 Conditional Entropy Interpretation of SCE Score
Note that the above provides our definition of the Schur-Complement-Entropy for the image part $\mathrm{SCE_{I}}$ and the text part $\mathrm{SCE_{T}}$ as follows:
$\mathrm{SCE}_{I}(x_{1},\ldots,x_{n})\,=$ $\,\sum_{j=1}^{d}\lambda^{(\Lambda_{I})}_{j}\log\frac{\mathrm{Tr}(\Lambda_{I})}{\lambda^{(\Lambda_{I})}_{j}}$ (9)
$\mathrm{SCE}_{T}(x_{1},\ldots,x_{n})\,:=$ $\,\sum_{j=1}^{d}\lambda^{(\Lambda_{T})}_{j}\log\frac{\mathrm{Tr}(\Lambda_{T})}{\lambda^{(\Lambda_{T})}_{j}}$ (10)
@@
Appendix B Additional Individual Image Decomposition Results via SC-Based Method
@@
Results for synthetic images generated using SDXL are shown in Figures 12 and 13. Corresponding results for DALL-E 3 are presented in Figures 14 and 15. Notably, the Schur Complement-based decomposition successfully isolates and removes image directions corresponding to a text condition that describes the concept to be excluded.
@@
Appendix C Additional Results on Diversity Evaluation
To further validate the findings presented in the main text, we conducted a similar experiment (Figure 5) using the Cosine Similarity Kernel. The results confirm that the diversity trends observed in the main text persist under a finite-dimensional kernel. Figures 16 and 17 illustrate the variation in SCE diversity when conditioned on different text prompts.
@@ Appendix C Additional Results on Diversity Evaluation
Moreover, we evaluate the diversity of typographically attacked ImageNet samples in Figure 21. Specifically, we overlay the text "cassette player" onto images from 10 different ImageNet classes and measure diversity as the number of distinct classes increases. The presence of overlaid text diverts CLIP’s sensitivity away from image content, causing it to encode the direction indicated by the text instead.
To illustrate this effect, we visualize salience maps of CLIP embeddings given a prompt referring to an object behind the overlayed text. The results show that CLIP is highly sensitive to centrally placed text, even when it is unrelated to the prompted object. However, applying SC-based decomposition mitigates this bias. This correction is reflected in the diversity plots, where $SCE_{I}$ increases rapidly as the number of ImageNet classes grows, whereas Vendi, Coverage, and Recall exhibit much weaker correlations with class diversity.
Figure 16: Plots by cancelling out ’cat’ and specific cat breed prompts (Cosine Sim Kernel)
@@
Figure 18: Evaluated SCE and Vendi scores with Gaussian Kernel on different animals with objects.
@@
Figure 21: Evaluated SCE, Vendi, Recall and Coverage scores with Gaussian Kernel on ImageNet with overlayed text.
Appendix D Results with Naive Text Embedding Subtraction without considering the adjust matrix $\Gamma^{*}$
In the given task setting, it may seem intuitive to assume that the difference between $\Phi_{I}$ and $\Phi_{T}$ would yield a similar outcome as the SC-based decomposition. To test this hypothesis, we compared the SC-based decomposition with a "naive" method, defined as $\Phi_{I}-\Phi_{T}$. In this naive approach, the learned correction matrix $\Gamma^{*}$ is replaced with an identity matrix, effectively omitting its computation. Our experiments reveal that such a decomposition usually fails to achieve the desired results and often leads to a loss of coherent directionality in the embedding space.
To evaluate the performance of the naive embedding subtraction method, we used the typographic attack dataset, which consists of 10 ImageNet classes where misleading text is overlaid on the images. We measured classification accuracy before and after decomposition. Figure 23 shows the distribution of classifications for images engraved with the text "cassette player." CLIP’s classification is heavily biased towards "cassette player," despite the underlying images belonging to a different class. After decomposition, the naive method removes the direction corresponding to the text but results in a skew towards "french horn," even though the image distribution is uniform across all 10 classes. In contrast, the SC-based decomposition corrects the embeddings, making them sensitive to the underlying images rather than the engraved text.

### v3

@@ 5 Numerical Results
We evaluated our proposed Schur Complement-based approach for CLIP decomposition across various scenarios using both real and synthetic datasets. Our experimental results are reported for the standard Cosine Similarity and Gaussian kernel functions. For the Gaussian kernel function, we use random Fourier features [46] with a random Fourier dimension $r=2000$ to embed the kernel using a finite feature dimension. In the diversity evaluation, we report our defined Scendi scores as well as the Vendi score [9], RKE[20], Recall[28] and Coverage[39] for unconditional diversity assessment without taking the text data into account. Note that Recall and Coverage may only be performed in the presence of a valid representative reference dataset. The clustering experiments were performed using Kernel PCA with the Gaussian kernel, highlighting the top clusters that align with the top eigenvector directions in the data.
@@
Figure 6: Clusters of DALL-E 3 [40] generated images of animals with fruits. The yellow column shows Kernel-PCA (KPCA) clusters using CLIP; the green column shows KPCA clusters with the application the proposed Schur-Complement-based method to remove ”fruit” direction from CLIP embedding; and the blue column shows clusters with the ”animal” direction removed from CLIP embedding.
metric prompt setting 1 shape 2 shapes 3 shapes
Vendi 5.93 7.03 9.46
$\mathrm{Scendi}$ w shapes 2.94 2.95 2.96
w/o shapes 2.53 2.96 3.70
Figure 7: Diversity measurement on the CLEVR dataset using Vendi and Scendi metrics. Vendi stays constant for all prompts, while $\mathrm{Scendi}$ only increases when shapes aren’t specified, highlighting true image-driven diversity.
Figure 8: LLM countries with Qwen3-0.6B [57]
Measuring Text-to-Image Model Diversity. To highlight the strengths of Scendi-based diversity evaluation, we conducted numerical experiments comparing unconditional and conditional diversity assessments. Figure 3 presents an experiment with the FLUX.1-dev [29] model, where individuals are generated based on an ambiguous prompt. We fixed the text prompt distribution and conditioned the model to generate doctors of varying gender and age. The results show that when the model is restricted to producing only ”senior male doctors,” the $\mathrm{Scendi}$ score is the lowest. Conversely, when the model generates a broader range of age groups and genders, the diversity score increases. We also provide an interpolated surface plot illustrating the intermediate mixing results. A diversity level of 0 indicates a strong bias toward a single characteristic (e.g., all images depict males), whereas a score of 1 signifies an equal representation of all characteristics (e.g., an equal number of male and female images).
@@
Figure 5 illustrates the inverse scenario: the image distribution is fixed while the accompanying text prompts vary in their specificity. We generated four sets of ”junior male doctor” images with different seeds, while prompts were assigned four levels of detail—fully specified (age and gender), partially specified (age only or gender only), or unspecified (no details). We measured $\mathrm{Scendi}$, Vendi, RKE, Coverage and Recall scores across four scenarios. Reference set for R/C is ”doctor” class in IdenProf dataset [38]. Vendi, RKE, Recall and Coverage scores remain relatively unchanged, whereas $\mathrm{Scendi}$ distinguishes between these cases. When the description is incomplete, $\mathrm{Scendi}$ is lowest due to the strong presence of ”junior male doctors” across all images, indicating bias. However, when prompts explicitly state that all images should depict ”junior male doctors,” diversity increases and is evaluated based on auxiliary features such as background details rather than age or gender. When partial information is given, $\mathrm{Scendi}$ falls between the fully specified and unspecified cases. These results highlight the need for differentiation between model-driven and prompt-driven diversity.
To further quantify our findings, in Figure 4 we conducted an experiment on SDXL generated cat images of different breeds. We generated 1,000 images per class and compared $\mathrm{Scendi}$ with other diversity metrics. When prompts did not specify cat breeds, the diversity captured by $\mathrm{Scendi}$ aligns with unconditional Vendi score. Conversely, when breed information was included in the prompts, diversity mainly stemmed from the text prompts rather than the image generator, leading to a non-increasing $\mathrm{Scendi}$ score. In both scenarios, Vendi, Coverage, and Recall scores remained unchanged. The reference set for Recall and Coverage consisted of ImageNet [7] ’cat’ samples. Additional experiments on other cases are discussed in the Appendix. We evaluated our metric on a subset of CLEVR [24] images containing only one shape to benchmark Scendi in a fully controlled environment. Each image was paired with two prompts: one specifying the shape and one omitting it, while keeping object count and colors constant. In a diversity evaluation analogous to Figure 4, similar trend remains.
This work primarily focuses on text‑to‑image evaluation using CLIP; however, the Scendi score naturally extends to any uni‑modal evaluation setting, such as text‑to‑text, which has become extremely popular with LLMs. Figure 8 mirrors our two previous experimental setups, but here the dataset consists of country prompts paired with short, fact‑only responses generated by Deepseek‑V3 [6]. Notably, this experiment demonstrates that in a uni‑modal context the Scendi score applies directly and does not require CLIP or other cross-modal embeddings, any suitable embedding will suffice.
@@
CLIP Decomposition. We use the eigenspace of the Schur complement component to interpret the diversity in the generated images that is independent of the text prompt, revealing the unique elements introduced by the generation model in the images. The eigenspace-based interpretation follows the application of the Kernel PCA (KPCA) clustering method and leads to a visualization of the sample clusters due to the input text clusters and the clusters with a feature added by the model to the generated image. Figure 6 shows the resulting KPCA-detected clusters for DALL-E 3 [40] generated images in response to prompts of type ”[animal] eating a [fruit]” with 5 different animal and fruit names. The detected clusters of CLIP image embedding (leftmost case) shows clusters with mixed animals and fruits. On the other hand, by considering the prompt of only specifying the fruit, the Schur complement component’s KPCA clusters highlight the animals in each cluster (middle case), and similarly by considering the prompt of only specifying the fruit, the animals are captured by each KPCA cluster of the Schur complement matrix. The results show how decomposing and removing the influence of concepts present in the prompts leads to changes in the major clusters. Further decomposition results and analysis are provided in the Appendix.
@@
Typographic Attacks. CLIP’s sensitivity to text within images, as showed in [37], makes it vulnerable to typographic attacks, where overlaying misleading text on an image influences its classification [31]. To investigate this, we constructed a dataset of 10 ImageNet classes, each image engraved with a random class label. Figure 9 demonstrates CLIP’s susceptibility, showing that top eigenvector directions capture the engraved text rather than the actual image content. By conditioning on prompts like ”text reading ’cassette player’ on top of image” and removing this direction, we re-clustered images based on the corrected CLIP embedding. This adjustment allowed us to identify clusters based on the actual image content rather than the engraved text. Additional results are provided in the Appendix.
@@ 6 Conclusion
In this work, we proposed a Schur Complement-based approach to decompose the kernel covariance matrix of CLIP image embeddings into the sum of text-induced and model-induced kernel covariance components. We demonstrated the application of this Schur Complemental approach to define the Scendi score that evaluates the diversity of prompt-guided generated data in a prompt-aware fashion. This method extends the application of CLIP embeddings from relevance evaluation to assessing diversity in text-to-image generation models. Additionally, we applied our approach to modify CLIP image embeddings by modifying or canceling the influence of an input text. Our numerical results showcase the potential for canceling text effects and its applications to computer vision tasks. Future research directions include extending the Schur Complement-based approach to other generative AI models, such as text-to-video and language models. Additionally, leveraging this decomposition to fine-tune the CLIP model for enhanced understanding of diverse concepts represents another future direction. Finally, applying Scendi score for diversity guidance in diffusion models, similar to RKE guidance in [22] and Vendi guidance in [13], will be relevant for future studies.
Acknowledgements
This work is partially supported by a grant from the Research Grants Council of the Hong Kong Special Administrative Region, China, Project 14209920, and is partially supported by CUHK Direct Research Grants with CUHK Project No. 4055164 and 4937054. Also, the authors would like to thank the anonymous reviewers and meta-reviewer for their constructive feedback and useful suggestions.
@@ A.1 Proof of Proposition 1
A.2 Conditional Entropy Interpretation of Scendi Score
As discussed in the main text, in Equation (6), both image component $\Lambda_{I}$ and text component $\Lambda_{T}$ are PSD matrices with unit trace. Furthermore, we have
@@ A.2 Conditional Entropy Interpretation of Scendi Score
Note that the above provides our definition of the Schur-Complement-Entropy for the image part $\mathrm{Scendi_{I}}$ and the text part $\mathrm{Scendi_{T}}$ as follows:
$\mathrm{Scendi}_{I}(x_{1},\ldots,x_{n})\,=$ $\,\sum_{j=1}^{d}\lambda^{(\Lambda_{I})}_{j}\log\frac{\mathrm{Tr}(\Lambda_{I})}{\lambda^{(\Lambda_{I})}_{j}}$ (8)
$\mathrm{Scendi}_{T}(x_{1},\ldots,x_{n})\,:=$ $\,\sum_{j=1}^{d}\lambda^{(\Lambda_{T})}_{j}\log\frac{\mathrm{Tr}(\Lambda_{T})}{\lambda^{(\Lambda_{T})}_{j}}$ (9)
@@
Appendix B Limitations
The Scendi framework is only compatible with cross‑modal embeddings, such as those produced by CLIP. When such embeddings are unavailable for a given data modality, evaluators cannot use Scendi to measure diversity. Extending Scendi to modalities without cross‑modal embeddings remains an open challenge and a promising direction for future work.
Appendix C Additional Individual Image Decomposition Results via SC-Based Method
@@
Results for synthetic images generated using SDXL are shown in Figures 10 and 11. Corresponding results for DALL-E 3 are presented in Figures 12 and 13. Notably, the Schur Complement-based decomposition successfully isolates and removes image directions corresponding to a text condition that describes the concept to be excluded.
@@
Appendix D Additional Results on Diversity Evaluation
To further validate the findings presented in the main text, we conducted a similar experiment (Figure 4) using the Cosine Similarity Kernel. The results confirm that the diversity trends observed in the main text persist under a finite-dimensional kernel. Figures 15 and 16 illustrate the variation in Scendi diversity when conditioned on different text prompts.
@@ Appendix D Additional Results on Diversity Evaluation
Moreover, we evaluate the diversity of typographically attacked ImageNet samples in Figure 21. Specifically, we overlay the text ”cassette player” onto images from 10 different ImageNet classes and measure diversity as the number of distinct classes increases. The presence of overlaid text diverts CLIP’s sensitivity away from image content, causing it to encode the direction indicated by the text instead.
To illustrate this effect, we visualize salience maps of CLIP embeddings given a prompt referring to an object behind the overlayed text. The results show that CLIP is highly sensitive to centrally placed text, even when it is unrelated to the prompted object. However, applying SC-based decomposition mitigates this bias. This correction is reflected in the diversity plots, where $Scendi_{I}$ increases rapidly as the number of ImageNet classes grows, whereas Vendi, Coverage, and Recall exhibit much weaker correlations with class diversity.
Figure 15: Plots by cancelling out ’cat’ and specific cat breed prompts (Cosine Similarity Kernel)
@@
Figure 17: Evaluated Scendi and Vendi scores with Gaussian Kernel on different animals with objects.
Figure 18: Classification distribution before and after CLIP correction on SDXL [44] generated images of animals with objects in the background
@@
Figure 21: Evaluated Scendi, Vendi, Recall and Coverage scores with Gaussian Kernel on ImageNet with overlayed text.
Appendix E Additional Experiments on the Image Captioning Task
In the main text, we discussed SC-based decomposition for text-to-image models and demonstrated how images can be decomposed given text prompts. Here, we show that the reverse process is also possible. Specifically, we explored decomposing captions based on their corresponding images.
The experimental setup mirrors that of Figure 18, with the key difference being that, instead of generating images for text prompts, we generated captions for the corresponding images. For this task, we used gpt4o-mini as the captioning model. Figure 19 illustrates the experimental setup.
We selected images closely aligned with the concept we aimed to remove. For instance, to eliminate the ”cat” direction in text, we used an image of a cat against a white background to better isolate the concept. After applying corrections for ”animals” or ”objects” in the text prompt, we observed successful decomposition, as reflected in the second column: the corrected CLIP embedding is no longer sensitive to the removed concept.
These findings highlight the versatility of the Scendi method, demonstrating its applicability across a wide range of tasks that rely on a shared embedding space.
Appendix F Robustness of Scendi
We note that the robustness of the Scendi framework depends on the choice of underlying embedding. To address limitations in CLIP, several alternatives have been introduced, such as FairCLIP [36]. Because Scendi is compatible with any cross‑modal embedding, we evaluated diversity using three additional variants: OpenCLIP [19], FairCLIP, and BLIP2 [33] on the SDXL generated cat breed dataset. The results appear in Figure 22. We observe that Scendi preserves the qualities of a diversity metric that increases as we introduce more breeds into the data pool.
Figure 22: Figure 5’s Scendi evaluation of different embeddings: OpenCLIP (left), Robust FAIR-CLIP (middle), BLIP2 (right)
Appendix G Results with Naive Text Embedding Subtraction without considering the adjustment matrix $\Gamma^{*}$
In the given task setting, it may seem intuitive to assume that the difference between $\Phi_{I}$ and $\Phi_{T}$ would yield a similar outcome as the SC-based decomposition. To test this hypothesis, we compared the SC-based decomposition with a ”naive” method, defined as $\Phi_{I}-\Phi_{T}$. In this naive approach, the learned correction matrix $\Gamma^{*}$ is replaced with an identity matrix, effectively omitting its computation. Our experiments reveal that such a decomposition usually fails to achieve the desired results and often leads to a loss of coherent directionality in the embedding space.
To evaluate the performance of the naive embedding subtraction method, we used the typographic attack dataset, which consists of 10 ImageNet classes where misleading text is overlaid on the images. We measured classification accuracy before and after decomposition. Figure 24 shows the distribution of classifications for images engraved with the text ”cassette player.” CLIP’s classification is heavily biased towards ”cassette player,” despite the underlying images belonging to a different class. After decomposition, the naive method removes the direction corresponding to the text but results in a skew towards ”french horn,” even though the image distribution is uniform across all 10 classes. In contrast, the SC-based decomposition corrects the embeddings, making them sensitive to the underlying images rather than the engraved text.